# MV-VTON: Multi-View Virtual Try-On with Diffusion Models

### v2

@@ Appendix B MORE QUALITATIVE RESULTS
Results on Multi-View VTON. In this section, we present more qualitative results on the MVG dataset. Specifically, as shown in Figure D, we show multiple groups of try-on results for the same person under different views, using the proposed method. In Figure D, the first column displays frontal-view and back-view garments, the second to fourth columns depict persons from different views, while the fifth to seventh columns showcase the corresponding try-on results. As can be seen, our method can generate realistic dressing-up results of the multi-view person from the given two views of clothing. Furthermore, our method can retain the details well on the original clothing (e.g., the buttons in the fifth row) and generate high-fidelity try-on images even under occlusion (e.g., hair occlusion in the second row). In conclusion, the proposed method exhibits outstanding performance on multi-view virtual try-on task.
Figure A. Visualization of the person and corresponding poses. We select one of garments based on the relative positions of left and right arms in the skeleton image when performing hard-selection on the multi-view virtual try-on task.
[…]
[Figure 1. Fully described in the text.]
Figure B. Examples of human parsing maps and dense pose in our dataset. The parsing maps can be used to synthesize cloth-agnostic person images.
[…]
[Figure 2. Fully described in the text.]
@@
Results on Frontal-View VTON. In this section, we show more visual comparison results on VITON-HD (Choi et al., 2021) dataset and DressCode (Morelli et al., 2022) dataset. Previous works include Paint By Example (Yang et al., 2023), PF-AFN (Ge et al., 2021b), GP-VTON (Xie et al., 2023), LaDI-VTON (Morelli et al., 2023), DCI-VTON (Gou et al., 2023) and StableVITON (Kim et al., 2023). The results are shown in Figure E. In the first and second row of Figure E, it can be seen that our method better preserves the shape of the original clothing (e.g., the cuff in the second row), compared to the previous methods. In addition, our method outperforms previous methods in preserving high-frequency details, such as patterns on clothing in the fourth and sixth rows. Moreover, in contrast to previous methods, MV-VTON is not constrained by specific types of clothing and can achieve highly realistic effects across a wide range of garment styles (e.g., the garment in the third row and the collar in the eighth row). In summary, our method also has superiority on frontal-view virtual try-on task.
@@
High Resolution Results on Frontal-View VTON. Following the previous works (Kim et al., 2023; Morelli et al., 2023; Gou et al., 2023), the above results are all obtained at 512$\times$384 resolution. In this section, we present more results at 1024$\times$768 resolution on VITON-HD (Choi et al., 2021) and DressCode (Morelli et al., 2022) datasets, as shown in the Figure F. Specifically, we utilize the model trained at 512$\times$384 resolution to directly test at 1024$\times$768 resolution. Despite the difference in resolutions between training and testing, our method can also produce high-fidelity try-on results. For instance, the generated images can preserve both the intricate patterns and text adorning the clothing (in the first row) while also effectively maintaining their original shapes (in the last row).
@@ Appendix C LIMITATIONS
Despite outperforming previous methods on both frontal-view and multi-view virtual try-on tasks, our method does not perform well in all cases. Figure C displays some unsatisfactory try-on results. As can be seen, although our method can preserve the shape and texture of original clothing (e.g., the ’DIESEL’ text in the first row), it is difficult for it to fully preserve some smaller or more complex details (e.g., the parts circled in red). The reason for this phenomenon may be that these details are easily lost when inpainting in latent space. We will try to solve this issue in future work.
Figure C. Visualization of bad cases on VITON-HD dataset.
[…]
[Figure 3. Fully described in the text.]
Figure D. Qualitative results on multi-view virtual try-on task.
[…]
[Figure 4. Fully described in the text.]
Figure E. Qualitative comparison on frontal-view virtual try-on task.
[…]
[Figure 5. Fully described in the text.]
Figure F. Qualitative results of 1024$\times$768 resolution on frontal-view virtual try-on task.
[…]
[Figure 6. Fully described in the text.]

### v4

@@ Appendix B MORE QUALITATIVE RESULTS
Comparison Results on Multi-View VTON. In this section, we present more comparison results on the MVG dataset. Specifically, in the first row of Figure C, due to the lack of adaptive selection of clothes, previous methods have difficulty in generating hoods of the original cloth. Moreover, in the second and third rows, previous methods often struggle to maintain fidelity to the original garments. In contrast, our method effectively addresses the aforementioned problems and generates high-fidelity results.
Figure C: Comparison results on multi-view virtual try-on task.
@@
Multi-View Results on Multi-View VTON. In this section, we present more multi-view results on the MVG dataset. Specifically, as shown in Figure E, we show multiple groups of try-on results for the same person under different views, using the proposed method. In Figure E, the first column displays frontal-view and back-view garments, the second to fourth columns depict persons from different views, while the fifth to seventh columns showcase the corresponding try-on results. As can be seen, our method can generate realistic dressing-up results of the multi-view person from the given two views of clothing. Furthermore, our method can retain the details well on the original clothing (e.g., the buttons in the fifth row) and generate high-fidelity try-on images even under occlusion (e.g., hair occlusion in the second row). In conclusion, the proposed method exhibits outstanding performance on multi-view virtual try-on task.
Complex Human Pose Results on Frontal-View VTON. In this section, we provide more VTON results under complex human pose conditions in Figure F. It can be seen that our method can also generate high-fidelity try-on results even when the target person has a more complex pose.
@@
Comparison Results on Frontal-View VTON. In this section, we show more visual comparison results on VITON-HD (Choi et al. 2021) dataset and DressCode (Morelli et al. 2022) dataset. The previous works include Paint By Example (Yang et al. 2023), PF-AFN (Ge et al. 2021b), GP-VTON (Xie et al. 2023), LaDI-VTON (Morelli et al. 2023), DCI-VTON (Gou et al. 2023) and StableVITON (Kim et al. 2023). The results are shown in Figure G. In the first and second row of Figure G, it can be seen that our method better preserves the shape of the original clothing (e.g., the cuff in the second row), compared to the previous methods. In addition, our method outperforms previous methods in preserving high-frequency details, such as patterns on clothing in the fourth and sixth rows. Moreover, in contrast to previous methods, MV-VTON is not constrained by specific types of clothing and can achieve highly realistic effects across a wide range of garment styles (e.g., the garment in the third row and the collar in the eighth row). In summary, our method also has superiority on frontal-view virtual try-on task.
@@
High Resolution Results on Frontal-View VTON. In this section, we present more results at 1024$\times$768 resolution on VITON-HD (Choi et al. 2021) and DressCode (Morelli et al. 2022) datasets, as shown in the Figure H. Specifically, we utilize the model trained at 512$\times$384 resolution to directly test at 1024$\times$768 resolution. Despite the difference in resolutions between training and testing, our method can also produce high-fidelity try-on results. For instance, the generated images can preserve both the intricate patterns and text adorning the clothing (in the first row) while also effectively maintaining their original shapes (in the last row).
Figure D: Visualization of bad cases on VITON-HD dataset.
@@ Appendix C LIMITATIONS
Despite outperforming previous methods on both frontal-view and multi-view virtual try-on tasks, our method does not perform well in all cases. Figure D displays some unsatisfactory try-on results. As can be seen, although our method can preserve the shape and texture of original clothing (e.g., the ’DIESEL’ text in the first row), it is difficult for it to fully preserve some smaller or more complex details (e.g., the parts circled in red). The reason for this phenomenon may be that these details are easily lost when inpainting in latent space. We will try to solve this issue in future work.
Figure E: Multi-view results on multi-view virtual try-on task.
Figure F: Complex human pose results on frontal-view virtual try-on task.
Figure G: Qualitative comparisons on frontal-view virtual try-on task.
Figure H: Qualitative results of 1024$\times$768 resolution on frontal-view virtual try-on task.